From mar@roe.ac.uk Mon Jul 29 14:56:58 2013 Date: Mon, 29 Jul 2013 14:56:50 +0100 From: Mike Read To: VDFS Science Archive Team -- Clive Davenhall , Eckhard Sutorius , Mark Holliman , Nicholas Cross , Nigel Hambly , Rob Blake , Ross Collins , Stelios Voutsinas Cc: CCs for VDFS -- Alastair Edge , Andrew Lawrence , Bob Mann , Dante Minniti , James Dunlop , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Maria-Rosa Cioni , Matt Jarvis , Mike Irwin , Nigel Metcalfe , Norman Gray , Phil Lucas , Richard McMahon at IOA , Simon Dye , Stephen Warren , Tom Kerr , Tom Shanks Subject: WFAU VDFS Science Archives progress meeting minutes, 26 July 2013 Present: MAR, RSC, ETWS, ACD, NJC Apologies: MSH, RGM, RPB, NCH, STV, KTN, AL, JPE DoNM: 10am Friday August 9th 2013 in the Villas meeting room NCH(?) in the chair (room is booked). Actions discharged this time: ----------------------------- None. Actions partly discharged but continuing: ----------------------------------------- ACTION: MAR to have a go at releasing other VISTA PI DBs (e.g. for JPE) Continues: MAR did not know if the release of 089.D-0113 had yet been successful as it had been causing problems, the release of JPE's project would follow. Actions carried forward from 12/07/2013 meeting ----------------------------------------------- ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues: this one still has low priority. ACTION: RPB to investigate maintaining consistency between WFAU and ESO releases. ACTION: All to think about items that could be included in the WFAU poster. Specific points and new actions: -------------------------------- Project management: WFAU quarterly to be held Tuesday afternoon 27th August 2013. WFCAM, VISTA and VST updates: Nothing to report. Comments and issues arising from CASU minutes: No new minutes this time. Networking: Up and working WSA/VSA/OSA/GES Operations: ETWS reported: -- Transferred the released GES iDR2 data for Giraffe and UVES. Copying over the WG data as it appears at CASU. -- Found quite a lot of missing OSA filter files (flats, confs) that CASU provides via symbolic links. Since some hiccup with WFCAM data we always excluded symbolic linked files, but this has been corrected now. -- VSA v1.3 catalogue ingests are running very well, SV and DRYRUN data is already in with P85 in progress. ETWS also noted that new OSA science files had also appeared on various days, including fixed versions of previously unfixed catalogues. Mike Irwin had been emailed to see what had prompted the generation of these files. Following a response a decision would be made on how best to handle them. Hardware and Systems: Ramses13, the VVV public server, was currently down/offline. There had been problems with the hardware (JBOD) being used for the MonetDB tests. Survey Data Release: The VMC release was nearly ready. Steve Warren had been in contact asking that UKIDSS DR9 be made world public. ACTION: MAR and RPB to make UKIDSS DR9 world public. Software: ETWS reported that he had: -- Parallelised CU4 which helps us running it at least as quick as the ingests into the database. -- Updated CU1 to get the GES data transferred. -- Updated the monitor page creation code to get file counts from disks quicker. NJC reported further work on the list driven curation code. RSC and ACD noted that they'd been working on the GES code and schema in preparation for ingesting and preparing the next release. Although the latest batch of data was more homogeneous it still did not quite match the supplied template. Some things still needed clarification but all the data/READMEs were not yet in place. New GES atomic line data had also been provided but in a different format than expected. Non-survey Data Release: Nothing to report. Astrogrid deployment, VO & Data Analysis services: Nothing to report. Miscellaneous: AL had updated the main WFAU webpage. NJC noted that PyFITS was now part of Astropy (astropy.io.fits). -- Scanned by iCritical.